# Efficient Pyramid Channel Attention Network for Pathological Myopia Recognition with Pretraining-and-Finetuning

### v2

@@ Abstract
Pathological myopia (PM) is the leading ocular disease for impaired vision worldwide. Clinically, the characteristic of pathology distribution in PM is global-local on the fundus image, which plays a significant role in assisting clinicians in diagnosing PM. However, most existing deep neural networks focused on designing complex architectures but rarely explored the pathology distribution prior of PM. To tackle this issue, we propose an efficient pyramid channel attention (EPCA) module, which fully leverages the potential of the clinical pathology prior of PM with pyramid pooling and multi-scale context fusion. Then, we construct EPCA-Net for automatic PM recognition based on fundus images by stacking a sequence of EPCA modules. Moreover, motivated by the recent pretraining-and-finetuning paradigm, we attempt to adapt pre-trained natural image models for PM recognition by freezing them and treating the EPCA and other attention modules as adapters. In addition, we construct a PM recognition benchmark termed PM-fundus by collecting fundus images of PM from publicly available datasets. The comprehensive experiments demonstrate the superiority of our EPCA-Net over state-of-the-art methods in the PM recognition task. The results also show that our method based on the pretraining-and-finetuning paradigm achieves competitive performance through comparisons to part of previous methods based on traditional fine-tuning paradigm with fewer tunable parameters, which has the potential to leverage more natural image foundation models to address the PM recognition task in limited medical data regime.
@@ 1 Introduction
Myopia or nearsightedness is one of the commonest ocular diseases. It is estimated that approximately 50% of people will have myopia by 2050 holden2016global in the world. The clinical symptoms of myopia are associated with multiple retinal changes such as fundus tessellation, retinal detachment, posterior staphyloma, atrophy, maculopathy, and so on chen2012types . Pathological myopia (PM) is one of the common-seen myopia types as well as the leading cause of irreversible visual impairment or blindness ruiz2019myopic . PM can increase the risk of pathologic changes for other ocular diseases such as glaucoma, retinal detachment, myopic macular degeneration and cataracts ohno2016updates . Therefore, it is important to recognize PM in time so that PM patients can take appropriate interventions and treatments to improve their vision and quality of life.
@@
Figure 1: Tessellated fundus information (red), parapapillary atrophy (blue) and optic disc (green).
@@
Over the years, deep neural networks, especially convolutional neural networks (CNNs), have achieved significant progress in ophthalmic image analysis, e.g., classification, lesion location, and semantic segmentation. Zhang et al. zhang2022adaptive proposed an adaptive feature squeeze network to classify nuclear cataracts. Jiang et al. JIANG2023104830 applied the Inception-V3 and ResNet to detect diabetic retinopathy (DR). He et al. he2020cabnet proposed a category attention network to grade DR. As for the PM recognition task, Du et al. DU20211235 applied a CNN model to detect subtle lesion features of myopic maculopathy based on fundus images. Tan et al. TAN2021e317 applied the pre-trained ResNet to automatic PM recognition with the full-tuning paradigm. Himami et al. 9689744 also used the pre-trained DenseNet to recognize PM automatically. Ye et al. ye2021automatic utilized the ResNet to identify myopic maculopathy. Although existing methods have achieved promising results, they have rarely infused the pathology distribution prior of PM into deep neural network design, limiting their applications to real clinical scenarios. Specifically, based on the fundus image, tessellated fundus information is distributed throughout the entire fundus area, while parapapillary atrophy and optic disc information are located in the specific areas, as shown in Fig. 1. From the above analysis, we gain the following sights: (1) The pathology distribution characteristic of PM is global-local on the fundus images, which CNNs can not exploit well. (2) Most PM recognition works used private datasets, lacking corresponding available PM recognition datasets.
@@
To address the first problem, we propose an Efficient Pyramid Channel Attention (EPCA) module, which explicitly infuses global-local pathology distribution prior into CNN representations through the context feature recalibration form, as shown in Fig. 2(c). It is comprised of a pyramid pooling operator and a multi-scale context fusion operator. The pyramid pooling operator extracts multi-scale context features from different feature map regions, widely used in channel attention block design, e.g., spatial pyramid attention (SPA) guo2020spanet ; ma2021spatial . Here, it is worth noting that different scale context features have varying levels of significance, but existing channel attention methods have rarely exploited how to fuse them in an efficient manner. In this way, we propose four kinds of multi-scale context feature fusion criteria to fulfill and validate the EPCA module: Linear-SCFM (multi-scale context feature mapping), Dropout-SCFM, Hierarchical-SCFM, and Parallel-SCFM. Then, this paper combines the EPCA module with the residual block to form the Res-EPCA module. Finally, we build the EPCA-Net for PM recognition based on fundus images by stacking a sequence of Res-EPCA modules. As for the problem of lacking public PM recognition datasets, we construct a new PM dataset named PM-Fundus by collecting fundus images of PM from publicly available fundus datasets: Pathological Myopia (PALM)  55pk-8z03-19 , Ocular Disease Intelligent Recognition (ODIR) ODIR ; zhou2020benchmark , and Retinal Fundus Multi-Disease Image (RFMiD) pachade2021retinal , which will be released as the PM recognition benchmark.
@@
Moreover, motivated by the advent of large-scale pre-trained models in natural language processing (NLP) and computer vision radford2021learning ; yuan2021florence ; yang2023aim , this paper attempts to adapt pre-trained natural image models to automatic PM recognition based on the pretraining-and-finetuning paradigm houlsby2019parameter ; hu2021lora . To implement this paradigm, we freeze pre-trained natural image models and take the proposed EPCA module and other attention methods as the adapters.
@@
    We construct an efficient pyramid channel attention network (EPCA-Net) for PM recognition based on fundus images based on multiple Res-EPCA modules. Moreover, we construct a PM dataset termed PM-Fundus that can be used as a benchmark dataset for PM recognition tasks. Extensive experiments on three PM datasets and the LAG dataset demonstrate the superiority of EPCA-Net over state-of-the-art methods.
@@ 2.1 Automatic PM Recognition
Over the past years, researchers have spent much effort into developing various CAD tools for PM and myopia to reduce the burden of clinicians and achieved significant progress SEPTIARINI2018151 through traditional machine learning methods and deep learning methods. In the machine learning-based PM recognition filed. Liu et al. liu2010detection ; tan2009automatic developed a PM recognize system named PAMELA (Pathological Myopia Detection Through Peripapillary Atrophy) based on fundus images. In the proposed PAMELA system, they focused on extracting texture features and clinical image context information, and then applied the support machine vector (SVM) to classify PM. Zhang et al. 10.1371/journal.pone.0065736 used retinal fundus imaging data, demographic and clinical information and genotyping data together for automatic PM classification with multiple kernel learning (MKL) methods. Cheng 6298012 applied biologically inspired features (BIF) for automatic myopia detection with sparse transfer learning. Li et al. LI2020105090 proposed an automatic parapapillary atrophy (PPA) detection framework for children’s myopia prediction.
Recently, deep neural networks have gradually been used for PM and myopia classification due to their powerful feature representation learning capability. Patil et al. 10.1007/978-981-19-2126-1_13 proposed an ensemble deep neural network framework for automatic myopia detection by ensembling DenseNet201, Xception and InceptionV3 together. Dai et al. 9102787 proposed a multi-task deep neural network to detect normal and abnormal and simultaneously recognize pathological myopia and high myopia. Wang et al. wang2023efficacy used the EfficientNet to recognize different PM stages. Chen et al. chen2023fit proposed a feature interaction transformer network (FIT-Net) to grade PM on optical coherence tomography (OCT) images. Sun et al. sun2023deep embedded the coarse clinical prior into ResNet50 for automatic myopic maculopathy grading. Li et al. li2022automated proposed a dual-stream CNN to classify PM. He et al. 9363019 proposed a modality-specific attention network (MSA-Net) to recognize PM. Although existing methods have achieved good PM recognition results, they often ignored how to explore global-local pathology distribution prior of PM to further improve the performance, specifically for fundus images. Moreover, we also find that most ophthalmic image datasets for PM recognition are private, lacking public PM recognition datasets.
@@ 2.2 Attention Mechanism
In recent years, attention mechanisms have demonstrated to be a powerful means to improve the representational capability of deep neural networks in a variety of computer vision tasks, such as image classification, semantic segmentation and object detection. The current research directions of attention mechanisms can be divided into channel attention, spatial attention and a combination of them.
More related to this paper is the channel attention mechanism. One of the most representative channel attention methods is the squeeze-and-excitation (SE) block hu2018squeeze , which extracts averaged global context features from feature maps with global average pooling and then apply them to model long-range dependencies among channels. Zhang et al. ZHANG2022109109 proposed the clinical-awareness attention (CCA) module to extract clinical features with mixed pooling and construct dependencies among channels via channel fully-connected layer. Wang et al. wang2020eca proposed an efficient channel attention (ECA) module to build local channel interaction. Lee et al. lee2019srm proposed a style-based recalibration (SRM) module to explore the potential of style features in CNNs. He et al. 9363019 presented a multi-scale attention (MSA) module to build relationships among feature representations from different stages of deep neural networks. Guo et al. guo2020spanet ; ma2021spatial developed a spatial pyramid attention (SPA) to extract multi-scale context information with the spatial pyramid pooling (SPP) and construct relationships between channels with two fully connected layers. Particularly, MSA and SPA are similar to our EPCA in extracting multi-scale context features that contain global and local pathology information, but they lose sight of how to fuse different scale context features in an efficient manner by considering their relative significance. Hence, unlike previous efforts, we design four kinds of fusion criteria for multi-scale context fusion in the proposed EPCA module from different aspects.
@@ 2.3 Parameter-efficient Transfer Learning
With the advent of large pre-trained models in NLP and computer vision, researchers found that transferring the large pre-trained models to other learning tasks with traditional finetuning methods became infeasible radford2021learning ; bommasani2021opportunities ; yuan2021florence ; yang2023aim . To address this limitation, parameter-efficient transfer learning has been proposed to adapt large pre-trained models to new learning tasks with few trainable parameters by freezing the pre-trained models ding2023parameter ; zaken2021bitfit ; chen2022adaptformer . Specifically, the adapter is one of the most popular parameter-efficient techniques, which plugs lightweight learnable modules into pre-trained models houlsby2019parameter ; yang2023aim ; lin2022swinbert . Various adapters like CLIP-based adapters gao2021clip ; lin2022frozen and ViT adapters chen2022vision ; gao2022visual have been developed for various tasks of NLP and computer vision. However, most adapter-based methods focused on adapting pre-trained models to the same domain (e.g., natural image-to-natural image). In contrast, this paper attempts to adapt the pre-trained natural image models to the PM recognition task by taking our proposed EPCA and other attention modules as adapters.
@@ 3.1 The overview of EPCA-Net
The overall structure of our EPCA-Net, as shown in Fig. 2(a). Our EPCA-Net first takes a fundus image as the input, and a convolutional layer is applied to produce coarse feature maps. Next, We combine the EPCA module with the Residual module to form a Res-EPCA module, as shown in Fig. 2(b). Then, we apply multiple Res-EPCA modules to extract augmented feature maps by dynamically adjusting the relative significance of multi-scale context features. Finally, a global average pooling layer is applied to obtain global feature representations from feature maps, and a softmax function is used to generate the predicted class of each input fundus image, which is illustrated in Fig. 2(a).
@@ 3.2.2 Multi-scale Context Fusion
Refer to caption (a) Drop-EPCA
Refer to caption (b) Hier-EPCA
Refer to caption (c) Par-EPCA
@@ 3.3 Pretraining-and-finetuning paradigm
Recently, with the advent of large pre-trained models, researchers found it is not easy to transfer large pre-trained models to other learning tasks with the traditional finetuning paradigm; this is mainly because these models require massive tunable parameters and computational cost. To address this limitation, the pretraining-and-finetuning paradigm, which has gained popularity in NLP and computer vision, has been proposed. However, only some existing works have adapted the pre-trained natural image models to medical image analysis with few tunable parameters. This paper attempts to directly adapt the pre-trained natural image models to PM detection based on the pretraining-and-finetuning paradigm. As shown in Fig. 4 (right lower part), we freeze pre-trained natural image models (e.g., ResNet50) and take our proposed EPCA as the adapter, which can deduce situations by treating other attention modules as adapters. Moreover, to understand the difference between the pretraining-and-finetuning paradigm and the traditional finetuning paradigm, Fig. 4 (left lower part) presents an example of the traditional finetuning paradigm.
Figure 4: The comparison between traditional finetuning paradigm and pretraining-and-finetuning paradigm.

### v3

@@ Abstract
Pathological myopia (PM) is the leading ocular disease for impaired vision worldwide. Clinically, the characteristic of pathology distribution in PM is global-local on the fundus image, which plays a significant role in assisting clinicians in diagnosing PM. However, most existing deep neural networks focused on designing complex architectures but rarely explored the pathology distribution prior of PM. To tackle this issue, we propose an efficient pyramid channel attention (EPCA) module, which fully leverages the potential of the clinical pathology prior of PM with pyramid pooling and multi-scale context fusion. Then, we construct EPCA-Net for automatic PM recognition based on fundus images by stacking a sequence of EPCA modules. Moreover, motivated by the recent pretraining-and-finetuning paradigm, we attempt to adapt pre-trained natural image models for PM recognition by freezing them and treating the EPCA and other attention modules as adapters. In addition, we construct a PM recognition benchmark termed PM-fundus by collecting fundus images of PM from publicly available datasets. The comprehensive experiments demonstrate the superiority of EPCA-Net over state-of-the-art methods in the PM recognition task. For example, EPCA-Net achieves 97.56% accuracy and outperforms ViT by 2.85% accuracy on the PM-fundus dataset. The results also show that our method based on the pretraining-and-finetuning paradigm achieves competitive performance through comparisons to part of previous methods based on traditional fine-tuning paradigm with fewer tunable parameters, which has the potential to leverage more natural image foundation models to address the PM recognition task in limited medical data regime.
@@ 1 Introduction
Myopia or nearsightedness is one of the commonest ocular diseases. It is estimated that approximately 50% of people will have myopia by 2050 [1] in the world. The clinical symptoms of myopia are associated with multiple retinal changes such as fundus tessellation, retinal detachment, posterior staphyloma, atrophy, maculopathy, and so on [2]. Pathological myopia (PM) is one of the common-seen myopia types as well as the leading cause of irreversible visual impairment or blindness [3]. PM can increase the risk of pathologic changes for other ocular diseases such as glaucoma, retinal detachment, myopic macular degeneration and cataracts [4]. Therefore, it is important to recognize PM in time so that PM patients can take appropriate interventions and treatments to improve their vision and quality of life.
@@
Figure 1: Tessellated fundus information (red), parapapillary atrophy (blue) and optic disc (green) on the fundus image.
@@
Over the years, deep neural networks, especially convolutional neural networks (CNNs), have achieved significant progress in ophthalmic image analysis, e.g., classification, lesion location, and semantic segmentation. Zhang et al. [5] proposed an adaptive feature squeeze network to classify nuclear cataracts. Jiang et al. [6] applied the Inception-V3 and ResNet to detect diabetic retinopathy (DR). He et al. [7] proposed a category attention network to grade DR. Fu et al. [8] combined U-Net with probability bubble to segment optic disc accurately in abnormal fundus images. As for the PM recognition task, Du et al. [9] applied a CNN model to detect subtle lesion features of myopic maculopathy based on fundus images. Tan et al. [10] applied the pre-trained ResNet to automatic PM recognition with the full-tuning paradigm. Himami et al. [11] also used the pre-trained DenseNet to recognize PM automatically. Ye et al. [12] utilized the ResNet to identify myopic maculopathy. Although existing methods have achieved promising results, they have rarely infused the pathology distribution prior of PM into deep neural network design, limiting their applications to real clinical scenarios. Specifically, based on the fundus image, tessellated fundus information is distributed throughout the entire fundus area, while parapapillary atrophy and optic disc information are located in the specific areas, as shown in Fig. 1. From the above analysis, we gain the following sights: (1) The pathology distribution characteristic of PM is global-local on the fundus images, which CNNs can not exploit well. (2) Most PM recognition works used private datasets, lacking corresponding available PM recognition datasets.
@@
To address the first problem, we propose an Efficient Pyramid Channel Attention (EPCA) module, which explicitly infuses global-local pathology distribution prior into CNN representations through the context feature recalibration form, as shown in Fig. 2(c). It is comprised of a pyramid pooling operator and a multi-scale context fusion operator. The pyramid pooling operator extracts multi-scale context features from different feature map regions, widely used in channel attention block design, e.g., spatial pyramid attention (SPA) [13]. Here, it is worth noting that different scale context features have varying levels of significance, but existing channel attention methods have rarely exploited how to fuse them in an efficient manner. In this way, we propose four kinds of multi-scale context feature fusion criteria to fulfill and validate the EPCA module: Linear-SCFM (multi-scale context feature mapping), Dropout-SCFM, Hierarchical-SCFM, and Parallel-SCFM. Then, this paper combines the EPCA module with the residual block to form the Res-EPCA module. Finally, we build the EPCA-Net for PM recognition based on fundus images by stacking a sequence of Res-EPCA modules. As for the problem of lacking public PM recognition datasets, we construct a new PM dataset named PM-Fundus by collecting fundus images of PM from publicly available fundus datasets: Pathological Myopia (PALM)  [14], Ocular Disease Intelligent Recognition (ODIR) [15, 16], and Retinal Fundus Multi-Disease Image (RFMiD) [17], which will be released as the PM recognition benchmark.
@@
Moreover, motivated by the advent of large-scale pre-trained models in natural language processing (NLP) and computer vision [18, 19, 20], this paper attempts to adapt pre-trained natural image models to automatic PM recognition based on the pretraining-and-finetuning paradigm [21, 22]. To implement this paradigm, we freeze pre-trained natural image models and take the proposed EPCA module and other attention methods as the adapters.
@@
    We construct an efficient pyramid channel attention network (EPCA-Net) for PM recognition based on fundus images by stacking multiple Res-EPCA modules. Moreover, we construct a PM dataset termed PM-Fundus that can be used as a benchmark dataset for PM recognition tasks. Extensive experiments on three PM datasets and the LAG dataset demonstrate the superiority of EPCA-Net over state-of-the-art methods.
@@ 2.1 Automatic PM Recognition
Over the past years, researchers have spent much effort into developing various CAD tools for PM and myopia to reduce the burden of clinicians and achieved significant progress [23] through traditional machine learning methods and deep learning methods. In the machine learning-based PM recognition filed. Liu et al. [24, 25] developed a PM recognize system named PAMELA (Pathological Myopia Detection Through Peripapillary Atrophy) based on fundus images. In the proposed PAMELA system, they focused on extracting texture features and clinical image context information, and then applied the support machine vector (SVM) to classify PM. Zhang et al. [26] used retinal fundus imaging data, demographic and clinical information and genotyping data together for automatic PM classification with multiple kernel learning (MKL) methods. Cheng [27] applied biologically inspired features (BIF) for automatic myopia detection with sparse transfer learning. Li et al. [28] proposed an automatic parapapillary atrophy (PPA) detection framework for children’s myopia prediction.
Recently, deep neural networks have gradually been used for PM and myopia classification due to their powerful feature representation learning capability. Patil et al. [29] proposed an ensemble deep neural network framework for automatic myopia detection by ensembling DenseNet201, Xception and InceptionV3 together. Dai et al. [30] proposed a multi-task deep neural network to detect normal and abnormal and simultaneously recognize pathological myopia and high myopia. Wang et al. [31] used the EfficientNet to recognize different PM stages. Chen et al. [32] proposed a feature interaction transformer network (FIT-Net) to grade PM on optical coherence tomography (OCT) images. Sun et al. [33] embedded the coarse clinical prior into ResNet50 for automatic myopic maculopathy grading. Li et al. [34] proposed a dual-stream CNN to classify PM. He et al. [35] proposed a modality-specific attention network (MSA-Net) to recognize PM. Although existing methods have achieved good PM recognition results, they often ignored how to explore global-local pathology distribution prior of PM to further improve the performance, specifically for fundus images. Moreover, we also find that most ophthalmic image datasets for PM recognition are private, lacking public PM recognition datasets.
@@ 2.2 Attention Mechanism
In recent years, attention mechanisms have demonstrated to be a powerful means to improve the representational capability of deep neural networks in a variety of computer vision tasks, such as image classification, semantic segmentation and object detection [36, 37, 38, 39, 13, 40]. The current research directions of attention mechanisms can be divided into channel attention, spatial attention and a combination of them.
More related to this paper is the channel attention mechanism. One of the most representative channel attention methods is the squeeze-and-excitation (SE) block [36], which extracts averaged global context features from feature maps with global average pooling and then apply them to model long-range dependencies among channels. Zhang et al. [37] proposed the clinical-awareness attention (CCA) module to extract clinical features with mixed pooling and construct dependencies among channels via channel fully-connected layer. Wang et al. [39] proposed an efficient channel attention (ECA) module to build local channel interaction. Fu et al. [38] developed a diabetic macular edema (DME) grading architecture by combing channel attention block with disease attention module. Lee et al. [41] proposed a style-based recalibration (SRM) module to explore the potential of style features in CNNs. He et al. [35] presented a multi-scale attention (MSA) module to build relationships among feature representations from different stages of deep neural networks. Guo et al. [13] developed a spatial pyramid attention (SPA) to extract multi-scale context information with the spatial pyramid pooling (SPP) and construct relationships between channels with two fully connected layers. Particularly, MSA and SPA are similar to our EPCA in extracting multi-scale context features that contain global and local pathology information, but they lose sight of how to fuse different scale context features in an efficient manner by considering their relative significance. Hence, unlike previous efforts, we design four kinds of fusion criteria for multi-scale context fusion in the proposed EPCA module from different aspects.
@@ 2.3 Parameter-efficient Transfer Learning
With the advent of large pre-trained models in NLP and computer vision, researchers found that transferring the large pre-trained models to other learning tasks with traditional finetuning methods became infeasible [18, 19, 20]. To address this limitation, parameter-efficient transfer learning has been proposed to adapt large pre-trained models to new learning tasks with few trainable parameters by freezing the pre-trained models [42, 43, 44]. Specifically, the adapter is one of the most popular parameter-efficient techniques, which plugs lightweight learnable modules into pre-trained models [21, 20, 45]. Various adapters like CLIP-based adapters [46, 47] and ViT adapters [48, 49] have been developed for various tasks of NLP and computer vision. However, most adapter-based methods focused on adapting pre-trained models to the same domain (e.g., natural image-to-natural image). In contrast, this paper attempts to adapt the pre-trained natural image models to the PM recognition task by taking our proposed EPCA and other attention modules as adapters.
@@ 3.1 The overview of EPCA-Net
The overall structure of our EPCA-Net, as shown in Fig. 2(a). Our EPCA-Net first takes a fundus image as the input, and a convolutional layer is applied to produce coarse feature maps. Next, We combine the EPCA module with the Residual module to form a Res-EPCA module, as shown in Fig. 2(b). Then, we apply multiple Res-EPCA modules to extract augmented feature maps by dynamically adjusting the relative significance of multi-scale context features. Finally, a global average pooling layer is applied to obtain global feature representations from feature maps, and a softmax function is used to generate the predicted class of each input fundus image, which is illustrated in Fig. 2(a). Table 1 provides the detailed architectures of EPCA-Nets based on ResNets, including input and output size of each layer.
@@
Table 1: The architectures of EPCA-Net18 and EPCA-Net50 for PM recognition based on fundus images
Output Size EPCA-Net18 EPCA-Net50
112$\times$112 7$\times$7, 64, stride=2
56$\times$56 3$\times$3, max pool, stride=2
56$\times$56 $\left[\begin{array}[]{cc}3\times 3,64\\ 3\times 3,64\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 64,64\end{array}\right]\times 2$ $\left[\begin{array}[]{cc}1\times 1,64\\ 3\times 3,64\\ 1\times 1,256\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 256,256\end{array}\right]\times 3$
28$\times$28 $\left[\begin{array}[]{cc}3\times 3,128\\ 3\times 3,128\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 128,128\end{array}\right]\times 2$ $\left[\begin{array}[]{cc}1\times 1,128\\ 3\times 3,128\\ 1\times 1,512\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 512,512\end{array}\right]\times 4$
14$\times$14 $\left[\begin{array}[]{cc}3\times 3,256\\ 3\times 3,256\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 256,256\end{array}\right]\times 2$ $\left[\begin{array}[]{cc}1\times 1,256\\ 3\times 3,256\\ 1\times 1,1024\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 1024,1024\end{array}\right]\times 6$
7$\times$7 $\left[\begin{array}[]{cc}3\times 3,512\\ 3\times 3,512\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 512,512\end{array}\right]\times 2$ $\left[\begin{array}[]{cc}1\times 1,512\\ 3\times 3,512\\ 1\times 1,2048\\ EPCA\left[S_{1}=1,S_{3}=9\right]\times 2048,2048\end{array}\right]\times 3$
1$\times$ 1 global average pool, fc, softmax
@@ 3.2.1 Pyramid Pooling
Figure 3: A simple implementation of pyramid pooling based on the $S_{1}$ and $S_{3}$.
@@ 3.2.2 Multi-scale Context Fusion
(a) Drop-EPCA
(b) Hier-EPCA
(c) Par-EPCA
@@ 3.3 Pretraining-and-finetuning paradigm
Recently, with the advent of large pre-trained models, researchers found it is not easy to transfer large pre-trained models to other learning tasks with the traditional finetuning paradigm. This is mainly because these models require massive tunable parameters and computational cost. The pretraining-and-finetuning paradigm has been proposed to address this limitation, which has gradually gained popularity in NLP and computer vision. However, only some existing works have adapted the pre-trained natural image models to medical image analysis with few tunable parameters. This paper attempts to directly adapt the pre-trained natural image models to PM detection based on the pretraining-and-finetuning paradigm. As shown in Fig. 5, we freeze pre-trained natural image models, e.g., ResNet50 (Left), and take our proposed EPCA as the adapter (Center), which can deduce situations by treating other attention modules as adapters. Moreover, to understand the difference between the pretraining-and-finetuning paradigm and the traditional finetuning paradigm, Fig. 5 (right) presents an example of the traditional finetuning paradigm, where pre-trained parameters of the residual block are finetuned during training.
Figure 5: TThe visual comparison between the traditional finetuning paradigm and the pretraining-and-finetuning paradigm by transferring the pre-trained natural image model of ResNet50 to tackle the PM recognition task.